OpenStudy (megannicole51):

Does the series converge? (-1)^(n-1)/(sqrt(3n-1)) n=1 to infinity

OpenStudy (anonymous):

this is an alternating series, all that is needed is that the terms go to zero

OpenStudy (megannicole51):

how did u know it was an alternating series?

OpenStudy (anonymous):

because of the \((-1)^{n-1}\) which is \(1\) if \(n\) is odd and \(-1\) if \(n\) is even

OpenStudy (megannicole51):

so then do i take the limit of it?

OpenStudy (anonymous):

no, you only need to check that the terms go to zero, which they do because, well because they do you have \(\frac{1}{\sqrt{3n-1}}\) the numerator is 1, and the denominator has an \(n\) in it, so the terms go to zero

OpenStudy (anonymous):

i mean i guess you have to check it, but really you just have to say it: \[\lim_{n\to \infty}\frac{1}{\sqrt{3n-1}}=0\]

OpenStudy (megannicole51):

okay awesome:)!!! thank you....now will it be that easy every time?

OpenStudy (anonymous):

i can imagine a harder example, but if the series alternates, and the terms go to zero, then the sum converges

OpenStudy (anonymous):

for example \[\sum\frac{1}{\ln(x)}\] certainly does not converge, but \[\sum \frac{(-1)^n}{\ln(n)}\] does

OpenStudy (megannicole51):

so do most alternating series converge?

OpenStudy (anonymous):

as long as the terms go to zero they do of course in your last example it would not, because the terms did not go to zero

OpenStudy (anonymous):

it is a much easier check that all the other stuff, integral test ratio test root test etc

OpenStudy (megannicole51):

yeah the ratio test is easy and the limit test

OpenStudy (anonymous):

btw you should recognize for a test or whatever that if you see \((-1)^n\) then the series alternates

OpenStudy (megannicole51):

okay awesome....the next problem however is telling me that it converges but is asking how many terms give a partial sum within .01 so do i just plug in terms and see which one will cross over .1 to tell me how many sums there are?

OpenStudy (anonymous):

hmm now i have to think that is certainly one way to do it by checking the numbers

OpenStudy (anonymous):

i think the error is no more that the absolute value of the last term

OpenStudy (anonymous):

what is the actual question? maybe we can do it

OpenStudy (megannicole51):

Find how many terms give a partial sum, Sn, within 0.01 of the sum, S, of the series. (-1)^(n-1)/(2n)! n=1 to infinity

OpenStudy (anonymous):

i am pretty sure that your job here is to find \(n\) so that \(\frac{1}{(2n!)}<.01\)

OpenStudy (anonymous):

you will need that many terms

OpenStudy (anonymous):

it won't be large that is for sure

OpenStudy (anonymous):

i think \(n=3\) will do it, since \(\frac{1}{6!}=\frac{1}{720}<.01\)

OpenStudy (megannicole51):

did u just pick 3 randomly?

OpenStudy (anonymous):

no i tries \(2\) first but \(\frac{1}{4!}=\frac{1}{24}\) is not small enough

OpenStudy (anonymous):

on the other hand we know \((2n)!\) gets real large real fast, so i figured to start small, i suppose you could say it was a guess

OpenStudy (megannicole51):

okay !!

OpenStudy (anonymous):

if it was \(\sum\frac{(-1)^n}{\sqrt{n}}\) we would have to solve \(\frac{1}{\sqrt{n}}<.01\) so in that case you would need a larger \(n\) like say \(n=1000\)

OpenStudy (anonymous):

actually \(n=10,000\)

OpenStudy (megannicole51):

where did the square root come from?